Points R, J, S, G are collinear. If RJ=SG and RJ, JS, and SG are integaers, what value is NOT possible for RG if JS= 15?
A: 18
B: 21
C: 25
D: 17

._____.__________._____.

the order of the points are R, J, S, G
: Points R, J, S, G are collinear. If RJ=SG and RJ, JS, and SG are integaers, what value is NOT possible for RG if JS= 15?
A: 18
B: 21
C: 25
D: 17

._____.__________._____.

the order of the points are R, J, S, G

Answer by gonzo(559)

order of points is RJSG (given)
RG is the whole line which is composed of the parts RJ, JS, and SG. (given)
JS = 15 (given)
RJ = SG (given)
RJ, JS, and SG are integers (given).
-----
RG = RJ + JS + SG (sum of the parts equals the whole).
RG = RJ + 15 + SG (substitution since JS = 15)
RG = 15 + 2*RJ (since RJ = SG, then we can susbsitute RJ for SG).
subtract 15 from both sides of equation.
RG - 15 = 2*RJ
divide both sides of equation by 2.
(RG - 15)/2 = RJ
-----
since RJ must an integer, then the equation on the left must also be an integer.
choices are:
18, 21, 25, 17
(18 - 13)/2 = 3/2 = 1.5 no good
(21 - 15)/2 = 6/2 = 3 ok
(25 - 15)/2 = 10/2 = 5 ok
(17 - 15)/2 = 2/2 = 1 ok
-----
answer is A (18).
RG cannot be 18 because (18-15)/2 equals 3/2 = 1.5 which is not an integer.
-----
answer is A.
JS cannot be equal to 1.5.